OpenStudy (anonymous):

find an equation of the ellipse with foci (√3,0) and (-√3,0) and sum of focal radii 4.

OpenStudy (anonymous):

sum of focal radiiii is equal to distance between the vertices 2a=4 a=2 distance b/w the focii=2ae 2ae=2sqrt3 ae=sqrt{3} e=sqrt{3}/2 b^2=4(1-3/4)=1 hence the eqn of ellipse is \[x ^{2}/4+y ^{2}/1=1\] hope its correct